One mole of a substance contains 6,022 × 10 per 23 protons, and since it is an electrically neutral system, it also contains an equal number of electrons. Suppose that we can properly extract protons from the original system and put them all together as one subsystem at a distance of one million meters from the electrons that remained at the original position and represent another subsystem. Determine the intensity of the force with which these two newly formed systems act on each other. (Solution: F=8,35 mN)


1
Expert's answer
2021-04-19T17:12:37-0400

The force can be found by Coulomb's law:


"F=k\frac{q_1q_2}{r^2}=k\frac{(ne)(ne)}{r^2}=k\frac{(ne)^2}{r^2},\\\space\\
F=9·10^9·\frac{(6.022·10^{23}·1.602·10^{-19})^2}{(10^6)^2}=83.78·10^6\text{ N}."